OpenStudy (anonymous):

Solve the inequality below for x: -4x-11≤13 A. x > 6 B. x < -6 C. x > -6 D. x < 6

OpenStudy (anonymous):

c is the correct answer, but really it should be \[x \le-6\]

OpenStudy (campbell_st):

add 11 to both sides -4x - 11 <= 13 + 11 + 11 _________________ -4x <= 24 divide by -4 and revserse the inequality x >= - 6

OpenStudy (anonymous):

oh wait my bad a. ≥6 b. x≤-6 c. x≥-6 d. x≤6

OpenStudy (anonymous):

sorry, it should be \[\ge\] used the wrong symbol